# Enhancing Whole Slide Pathology Foundation Models through Stain Normalization

### v1

@@ 4 Experiments
Table 1: Number of slides and patches per source in the pretraining data. EXAPath is trained with 285,153,903 patches
Data Source # WSIs # patches
TCGA 11,677 154,648,179
GTEx 23,118 130,505,724
Total 34,795 285,153,903
Figure 3: MPP distribution of the training data. (a) The MPP distribution of patches from the TCGA dataset. (b) The MPP distribution of patches from the GTEx dataset. (c) The overall MPP distribution of patches from both the TCGA and GTEx datasets.
Table 1 shows the number of whole slides and patches per data source used for pretraining. We first collected 11,677 Formalin-Fixed, Paraffin-Embedded (FFPE) whole slides from The Cancer Genome Atlas (TCGA) (Weinstein et al., 2013) and 23,118 FFPE whole slides from the Genotype-Tissue Expression (GTEx) project (Lonsdale et al., 2013). Both datasets consist of Hematoxylin and Eosin (H&E) stained WSIs. The TCGA dataset contains WSIs for various cancer types, while the GTEx dataset contains WSIs of normal tissues. Both the TCGA and GTEx datasets are publicly available. Following CLAM (Lu et al., 2021), we divide the WSIs into non-overlapping patches, focusing only on the tissue-containing regions. Given that each WSI has a different micron per pixel (MPP), we adjust the patch sizes so that they correspond to 0.5 MPP when resized to $256^{2}$. For instance, for WSIs with 0.25 MPP, we extract patches at $512^{2}$ dimensions, and for WSIs with 0.5 MPP, we extract patches at $256^{2}$ dimensions. However, for training the EXAPath model, the patches are not resized to achieve a uniform 0.5 MPP. Instead, we retain the original patch sizes during training because the DINO algorithm’s input image transform includes random resized cropping. Even if we resize the patches to 0.5 MPP, the model’s input would still undergo changes in MPP due to this random resized crop operation. Therefore, maintaining the originally extracted patch sizes allows us to preserve the maximum amount of information from each WSI while relying on DINO’s augmentations to introduce the necessary variability in scale and perspective. Figure 3(a) shows the distribution of MPP values for the patches extracted from the TCGA dataset, and Figure 3(b) shows the distribution of MPP values for the patches extracted from the GTEx dataset. Figure 3(c) presents the overall distribution of MPP values for patches extracted from both the TCGA and GTEx datasets. Most of the data used for training is concentrated around 0.25 MPP and 0.5 MPP. The total number of patches used for training EXAPath is 285,153,903.
@@
We train EXAPath using DINO (Caron et al., 2021) with a ViT-B model (Dosovitskiy et al., 2021), which has a patch size of 16, initializing the weights with those pretrained on the ImageNet dataset (Russakovsky et al., 2015). Training is conducted on 10 machines, each equipped with 8 A100 GPUs. We use a batch size of 5,120 and a learning rate of 0.005, training the model for 10 epochs. The learning rate is warmed up over the first 1,000 iterations. We employ bfloat16 precision training. The output layer is not fixed; instead, it is trained from scratch. Additionally, we set the number of local crops to 8. Other hyperparameters follow the settings from the original DINO paper for training the ViT-B model.
@@
We evaluate EXAPath on six publicly available patch-level datasets for classification tasks: PCAM (Veeling et al., 2018; Bejnordi et al., 2017), MHIST (Wei et al., 2021), CRC-100K (Kather et al., 2018), TIL Detection (Abousamra et al., 2022; Saltz et al., 2018b; Kaczmarzyk et al., ), MSI CRC (Kather, 2019), and MSI STAD (Kather, 2019).
@@
PCAM (Veeling et al., 2018; Bejnordi et al., 2017) consists of 327,680 color images, each of $96^{2}$ pixels, extracted from histopathologic scans of lymph node sections, and each annotated with a binary label indicating the presence of metastatic tissue. The dataset is divided into a training set of 262,144 examples and validation and test sets of 32,768 examples each, with no overlap between WSIs across splits. All splits maintain a 50/50 balance between positive and negative examples. A positive label indicates the presence of at least one pixel of tumor tissue within the central $32^{2}$ pixel region of a patch, while tumor tissue in the outer region does not affect the label. PCAM is derived from the Camelyon16 Challenge, which includes 400 H&E-stained WSIs of sentinel lymph node sections, digitized at 40x magnification (0.243 MPP) and undersampled to 10x for a larger field of view. The dataset follows the train/test split from Camelyon16, with 20% of the training WSIs held out for validation.
@@
MHIST (Wei et al., 2021) consists of 3,152 H&E-stained FFPE images, each $224^{2}$ pixels in size, of colorectal polyps, collected from the Department of Pathology and Laboratory Medicine at Dartmouth-Hitchcock Medical Center (DHMC). Each image is scanned at 40x magnification and resized to 8x magnification and is labeled based on the consensus of seven pathologists at DHMC, categorizing the type of colorectal polyps. The MHIST dataset focuses on a binary classification task, distinguishing between Hyperplastic Polyps (HP) and Sessile Serrated Adenomas (SSA). The dataset is divided into training and test sets, with the training set further split into training and validation sets using an 80:20 ratio. This results in 1,740 training samples, 435 validation samples, and 977 test samples.
@@
CRC-100K (Kather et al., 2018) consists of 100,000 non-overlapping image patches ($224^{2}$ pixels, 0.5 MPP) from H&E-stained histological images of human colorectal cancer (CRC) and normal tissue. The dataset includes both color-normalized images using Macenko’s method, and non-normalized images, with the latter exhibiting slight variations in staining intensity and color. We utilize the non-normalized images for our work, as we intend to apply Macenko normalization using our Macenko target image. Tissue classes include Adipose (ADI), Background (BACK), Debris (DEB), Lymphocytes (LYM), Mucus (MUC), Smooth Muscle (MUS), Normal Colon Mucosa (NORM), Cancer-associated Stroma (STR), and Colorectal Adenocarcinoma Epithelium (TUM). These patches are manually extracted from 86 H&E-stained FFPE samples from the NCT Biobank and the UMM pathology archive. An additional set, CRC-VAL-HE-7K, consists of 7,180 patches from 50 patients with CRC and serves as a validation set with no overlap with the original dataset. Since there is no separate test set provided, we use CRC-VAL-HE-7K as our test set. For training and validation, we split the original training set into an 80:20 ratio to create training and validation sets.
@@
TIL Detection (Abousamra et al., 2022; Saltz et al., 2018b; Kaczmarzyk et al., ) dataset consists of 304,097 H&E-stained images, each $100^{2}$ pixels at 0.5 MPP, extracted from FFPE sections covering 23 different cancer types from TCGA. The dataset is used for predicting TIL-positive images, where an image is considered positive if it contains more than two tumor-infiltrating lymphocytes (TILs). The dataset is divided into 209,221 training images, 38,601 validation images, and 56,275 test images. Each image is annotated as either TIL-positive (54,910 images) or TIL-negative (249,187 images), based on the presence of at least two TILs.
@@
MSI CRC and MSI STAD (Kather, 2019). For MSI CRC, the training set contains 93,408 image patches, and the test set comprises 99,904 image patches. For MSI STAD, the training set includes 100,570 image patches, and the test set contains 118,008 image patches. All images are $224^{2}$ pixels at 0.5 MPP and are derived from histological images of colorectal and gastric cancer patients in the TCGA cohort. The images are sourced from FFPE diagnostic slides. Preprocessing includes automatic tumor detection, resizing to $224^{2}$ pixels, and color normalization using Macenko’s method. Patients are categorized into ”MSS” (microsatellite stable) or ”MSIMUT” (microsatellite instable or highly mutated) groups, and the data is split into training (70%) and testing (30%) sets at the patient level, with the training set balanced by undersampling excess MSS tiles. We further split the training set into training and validation sets with an 80:20 ratio.
@@ 4.4 Downstream Task Training Details
We use linear evaluation, a common practice in previous SSL research for evaluation (Chen et al., 2020a; Zbontar et al., 2021; He et al., 2020). Training is performed using the SGD optimizer with a learning rate of 0.1, without weight decay, and a momentum of 0.9, with a batch size of 128 for 12,500 iterations. The images are first resized to $256^{2}$ pixels and then center-cropped to $224^{2}$ pixels. No additional data augmentation is applied. Macenko normalization is used in both the training and evaluation processes. Even datasets that are already Macenko-normalized are re-normalized using the same Macenko target images employed during EXAPath training.
@@ 4.5 Downstream Task Results
Table 2: Linear evaluation performance on six downstream tasks. EXAPath is compared with other models, including state-of-the-art pathology foundation models. Top-1 accuracy is shown. Values for models other than GigaPath are taken from the RudolfV paper
@@
EXAPath (ours) 35k 285M 86M 0.901 0.818 0.946 0.939 0.756 0.804 0.861
Table 2 presents the classification accuracy results obtained from the linear evaluation. EXAPath demonstrated competitive performance, achieving an average accuracy of 0.861 across the six downstream tasks. Notably, EXAPath achieves the highest performance among all models on the MSI CRC dataset, with an accuracy of 0.756, and on the MSI STAD dataset, with an accuracy of 0.804. Figure 1(a) compares the number of model parameters with the average accuracy across the six downstream tasks. Models positioned in the top left have fewer parameters and higher accuracy, indicating better efficiency. EXAPath exhibits parameter efficiency compared to other models. Figure 1(b) shows the average accuracy across six downstream tasks in relation to the number of WSIs used for model training. Models in the top left exhibit higher accuracy with fewer WSIs used for training. Training with stain-normalized images effectively mitigates the WSI-specific feature collapse problem, allowing for strong performance with fewer data and relatively smaller models.
@@ 5 Conclusion
In this paper, we have made a notable discovery during the visualization of features from patches extracted from models trained with self-supervised learning. We have termed this phenomenon ‘WSI-specific feature collapse’, where features from patches extracted from the same Whole Slide Image (WSI) tend to cluster together. This suggests inefficient learning, as it leads to the extraction of similar features when thousands of patches are derived from a single WSI. We hypothesized that this phenomenon arose from the substantial color variations between different WSIs. To address this issue, we utilized Macenko normalization on the training data to minimize color differences and developed EXAPath. As a result, EXAPath significantly reduced the WSI-specific feature collapse and achieved comparable or superior performance to existing models, even though those models use more data and parameters. However, despite the considerable improvement in mitigating WSI-specific feature collapse through EXAPath, this phenomenon appears to persist to some extent. This suggests the need for further research. Specifically, developing more effective and computationally efficient stain normalization techniques, suitable for integration into data loaders, could be a promising direction for future research. Additionally, exploring new learning methods or model architectures that can more effectively prevent feature collapse is also necessary. In conclusion, EXAPath represents significant progress in WSI analysis, laying the groundwork for more efficient and accurate pathological image analysis.

### v3

@@ 4 Experiments
Figure 3: MPP distribution of the training data. Most of the data used for training is concentrated around 0.25 MPP and 0.5 MPP.
We collected 34,795 Formalin-Fixed, Paraffin-Embedded (FFPE) Hematoxylin and Eosin (H&E) stained whole slide images (WSIs). Following CLAM [Lu et al., 2021], we divide the WSIs into non-overlapping patches, focusing only on the tissue-containing regions. Given that each WSI has a different micron per pixel (MPP), we adjust the patch sizes so that they correspond to 0.5 MPP when resized to $256^{2}$. For instance, for WSIs with 0.25 MPP, we extract patches at $512^{2}$ dimensions, and for WSIs with 0.5 MPP, we extract patches at $256^{2}$ dimensions. However, for training the EXAONEPath model, the patches are not resized to achieve a uniform 0.5 MPP. Instead, we retain the original patch sizes during training because the DINO algorithm’s input image transform includes random resized cropping. Even if we resize the patches to 0.5 MPP, the model’s input would still undergo changes in MPP due to this random resized crop operation. Therefore, maintaining the originally extracted patch sizes allows us to preserve the maximum amount of information from each WSI while relying on DINO’s augmentations to introduce the necessary variability in scale and perspective. Figure 3 presents the distribution of MPP values for patches extracted from the WSIs. Most of the data used for training is concentrated around 0.25 MPP and 0.5 MPP. The total number of patches used for training EXAONEPath is 285,153,903.
@@
We train EXAONEPath using DINO [Caron et al., 2021] with a ViT-B model [Dosovitskiy et al., 2021], which has a patch size of 16, initializing the weights with those pretrained on the ImageNet dataset [Russakovsky et al., 2015]. Training is conducted on 10 machines, each equipped with 8 A100 GPUs. We use a batch size of 5,120 and a learning rate of 0.005, training the model for 10 epochs. The learning rate is warmed up over the first 1,000 iterations. We employ bfloat16 precision training. The output layer is not fixed; instead, it is trained from scratch. Additionally, we set the number of local crops to 8. Other hyperparameters follow the settings from the original DINO paper for training the ViT-B model.
@@
We evaluate EXAONEPath on six publicly available patch-level datasets for classification tasks: PCAM [Veeling et al., 2018, Bejnordi et al., 2017], MHIST [Wei et al., 2021], CRC-100K [Kather et al., 2018], TIL Detection [Abousamra et al., 2022, Saltz et al., 2018b, Kaczmarzyk et al., ], MSI CRC [Kather, 2019], and MSI STAD [Kather, 2019].
@@
PCAM [Veeling et al., 2018, Bejnordi et al., 2017] consists of 327,680 color images, each of $96^{2}$ pixels, extracted from histopathologic scans of lymph node sections, and each annotated with a binary label indicating the presence of metastatic tissue. The dataset is divided into a training set of 262,144 examples and validation and test sets of 32,768 examples each, with no overlap between WSIs across splits. All splits maintain a 50/50 balance between positive and negative examples. A positive label indicates the presence of at least one pixel of tumor tissue within the central $32^{2}$ pixel region of a patch, while tumor tissue in the outer region does not affect the label. PCAM is derived from the Camelyon16 Challenge, which includes 400 H&E-stained WSIs of sentinel lymph node sections, digitized at 40x magnification (0.243 MPP) and undersampled to 10x for a larger field of view. The dataset follows the train/test split from Camelyon16, with 20% of the training WSIs held out for validation.
@@
MHIST [Wei et al., 2021] consists of 3,152 H&E-stained FFPE images, each $224^{2}$ pixels in size, of colorectal polyps, collected from the Department of Pathology and Laboratory Medicine at Dartmouth-Hitchcock Medical Center (DHMC). Each image is scanned at 40x magnification and resized to 8x magnification and is labeled based on the consensus of seven pathologists at DHMC, categorizing the type of colorectal polyps. The MHIST dataset focuses on a binary classification task, distinguishing between Hyperplastic Polyps (HP) and Sessile Serrated Adenomas (SSA). The dataset is divided into training and test sets, with the training set further split into training and validation sets using an 80:20 ratio. This results in 1,740 training samples, 435 validation samples, and 977 test samples.
@@
CRC-100K [Kather et al., 2018] consists of 100,000 non-overlapping image patches ($224^{2}$ pixels, 0.5 MPP) from H&E-stained histological images of human colorectal cancer (CRC) and normal tissue. The dataset includes both color-normalized images using Macenko’s method, and non-normalized images, with the latter exhibiting slight variations in staining intensity and color. We utilize the non-normalized images for our work, as we intend to apply Macenko normalization using our Macenko target image. Tissue classes include Adipose (ADI), Background (BACK), Debris (DEB), Lymphocytes (LYM), Mucus (MUC), Smooth Muscle (MUS), Normal Colon Mucosa (NORM), Cancer-associated Stroma (STR), and Colorectal Adenocarcinoma Epithelium (TUM). These patches are manually extracted from 86 H&E-stained FFPE samples from the NCT Biobank and the UMM pathology archive. An additional set, CRC-VAL-HE-7K, consists of 7,180 patches from 50 patients with CRC and serves as a validation set with no overlap with the original dataset. Since there is no separate test set provided, we use CRC-VAL-HE-7K as our test set. For training and validation, we split the original training set into an 80:20 ratio to create training and validation sets.
@@
TIL Detection [Abousamra et al., 2022, Saltz et al., 2018b, Kaczmarzyk et al., ] dataset consists of 304,097 H&E-stained images, each $100^{2}$ pixels at 0.5 MPP, extracted from FFPE sections covering 23 different cancer types from TCGA [Weinstein et al., 2013]. The dataset is used for predicting TIL-positive images, where an image is considered positive if it contains more than two tumor-infiltrating lymphocytes (TILs). The dataset is divided into 209,221 training images, 38,601 validation images, and 56,275 test images. Each image is annotated as either TIL-positive (54,910 images) or TIL-negative (249,187 images), based on the presence of at least two TILs.
@@
MSI CRC and MSI STAD [Kather, 2019]. For MSI CRC, the training set contains 93,408 image patches, and the test set comprises 99,904 image patches. For MSI STAD, the training set includes 100,570 image patches, and the test set contains 118,008 image patches. All images are $224^{2}$ pixels at 0.5 MPP and are derived from histological images of colorectal and gastric cancer patients in the TCGA cohort. The images are sourced from FFPE diagnostic slides. Preprocessing includes automatic tumor detection, resizing to $224^{2}$ pixels, and color normalization using Macenko’s method. Patients are categorized into "MSS" (microsatellite stable) or "MSIMUT" (microsatellite instable or highly mutated) groups, and the data is split into training (70%) and testing (30%) sets at the patient level, with the training set balanced by undersampling excess MSS tiles. We further split the training set into training and validation sets with an 80:20 ratio.
@@ 4.4 Downstream Task Training Details
We use linear evaluation, a common practice in previous SSL research for evaluation [Chen et al., 2020a, Zbontar et al., 2021, He et al., 2020]. Training is performed using the SGD optimizer with a learning rate of 0.1, without weight decay, and a momentum of 0.9, with a batch size of 128 for 12,500 iterations. The images are first resized to $256^{2}$ pixels and then center-cropped to $224^{2}$ pixels. No additional data augmentation is applied. Macenko normalization is used in both the training and evaluation processes. Even datasets that are already Macenko-normalized are re-normalized using the same Macenko target images employed during EXAONEPath training.
@@ 4.5 Downstream Task Results
Table 1: Linear evaluation performance on six downstream tasks. EXAONEPath is compared with other models, including state-of-the-art pathology foundation models. Top-1 accuracy is shown. Values for models other than GigaPath are taken from the RudolfV paper
@@
EXAONEPath (ours) 35k 285M 86M 0.901 0.818 0.946 0.939 0.756 0.804 0.861
Table 1 presents the classification accuracy results obtained from the linear evaluation. EXAONEPath demonstrated competitive performance, achieving an average accuracy of 0.861 across the six downstream tasks. Notably, EXAONEPath achieves the highest performance among all models on the MSI CRC dataset, with an accuracy of 0.756, and on the MSI STAD dataset, with an accuracy of 0.804. Figure 1(a) compares the number of model parameters with the average accuracy across the six downstream tasks. Models positioned in the top left have fewer parameters and higher accuracy, indicating better efficiency. EXAONEPath exhibits parameter efficiency compared to other models. Figure 1(b) shows the average accuracy across six downstream tasks in relation to the number of WSIs used for model training. Models in the top left exhibit higher accuracy with fewer WSIs used for training. Training with stain-normalized images effectively mitigates the WSI-specific feature collapse problem, allowing for strong performance with fewer data and relatively smaller models.
@@ 5 Conclusion
In this paper, we have made a notable discovery during the visualization of features from patches extracted from models trained with self-supervised learning. We have termed this phenomenon ‘WSI-specific feature collapse’, where features from patches extracted from the same Whole Slide Image (WSI) tend to cluster together. This suggests inefficient learning, as it leads to the extraction of similar features when thousands of patches are derived from a single WSI. We hypothesized that this phenomenon arose from the substantial color variations between different WSIs. To address this issue, we utilized Macenko normalization on the training data to minimize color differences and developed EXAONEPath. As a result, EXAONEPath significantly reduced the WSI-specific feature collapse and achieved comparable or superior performance to existing models, even though those models use more data and parameters. However, despite the considerable improvement in mitigating WSI-specific feature collapse through EXAONEPath, this phenomenon appears to persist to some extent. This suggests the need for further research. Specifically, developing more effective and computationally efficient stain normalization techniques, suitable for integration into data loaders, could be a promising direction for future research. Additionally, exploring new learning methods or model architectures that can more effectively prevent feature collapse is also necessary. In conclusion, EXAONEPath represents significant progress in WSI analysis, laying the groundwork for more efficient and accurate pathological image analysis.